A force of 50 kgf is applied to the smaller piston of a hydraulic machine. Neglecting the friction, find the force exerted on the large piston, the diameters of the pistons being 5 cm and 25 cm respectively.

Grade:9

1 Answers

Arun
6 years ago
student-name Harsh Vardhan asked in PhysicsA force of 50 kgf is applied to the smaller piston of a hydraulic machine.Neglecting friction,find the force exerted on the large piston, the diameters of the pistons being 5cm and 25 cm respectively. 1 Follow 0student-name Radhika Tandon answered this39 helpful votes in Physics, Class XII-Sciencehi...actually the pascals law says that pressure applied to any point in a liquid will be transmitted equally to all points in liquids..no matter the force and area can differ..so keeping this statement;let force and area of smaller piston be F1and A1..force and area of larger piston be F2and A2..radius of smaller piston=5/2=2.5cmradius of large piston=25/2since, pressure=force/areaso, F1/ A1=F2/ A2A1=pie(r)2=pie(2.5)2 cmA2=pie(r)2=pie (12.5)2 cm50/pie 2.5*2.5=F2/pie 12.5*12.550/2.5*2.5=F2/12.5*12.5F2=50*12.5*12.5/6.25=1250 kgfhence force on large piston =1250kgf